OpenStudy (anonymous):

What is true about the complex number -5 - 5i? Check all that apply. A. The modulus is 5. B. It lies in quadrant 3. C. It lies in quadrant 2. D. The modulus is 5 square root of 2.

OpenStudy (mertsj):

\[modulus = \sqrt{a^2+b^2}\]

OpenStudy (mertsj):

Left and down. What quadrant is that?

OpenStudy (anonymous):

3?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

So only B is the right answer

OpenStudy (mertsj):

nope

OpenStudy (jkristia):

thre is one more answer

OpenStudy (mertsj):

What's the modulus?

OpenStudy (mertsj):

modulus = square root of a^2+b^2

OpenStudy (anonymous):

OOOHHHH i see yeah, ur right

OpenStudy (jkristia):

no modulus is not 5

OpenStudy (amorfide):

3rd quadrant -5=x -5i=y (x,y) -5 on x axes -5 on y axis

OpenStudy (amorfide):

modulus is the square root of (-5²+-5²) which is root 50

OpenStudy (amorfide):

root 50 = 5root 2 B and D are correct

OpenStudy (amorfide):

sorry for so many posts, had to figure it out in my head since i ran out of paper

OpenStudy (jkristia):

I thought the purpose of this site is the help with math problem by giving pointers instead of just solving them.

OpenStudy (anonymous):

I already figured it out by following modulus = square root of a^2+b^2 its just clarification no need to be sensitive..

OpenStudy (amorfide):

okay -5²+-5²= root 50 root 50= root 25 x root 2 root 25=5 5 root 2

OpenStudy (amorfide):

if there is anything else not explained, let me know

OpenStudy (anonymous):

Thanks